# Ancora: Accurate Intrusion Recovery for Web Applications

### v2

@@ Abstract
Modern web application recovery presents a critical dilemma. Coarse-grained snapshot rollbacks cause unacceptable data loss for legitimate users. However, surgically removing an attack’s impact is hindered by a fundamental challenge in high-concurrency environments: it is difficult to attribute the resulting file, database modifications, etc., to a specific attack related request.
We present Ancora, a system for precise intrusion recovery in web applications without invasive instrumentation. Our approach first isolates the full sequence of syscalls triggered by a single malicious request. Based on this sequence, Ancora addresses file and database modifications separately. To trace file modifications, it builds a provenance graph that reveals all changes, even those made by exploit-spawned processes. To attribute database operations, a more difficult challenge due to connection pooling, Ancora introduces a novel technique based on a spatiotemporal anchor. This anchor uses the request’s network connection tuple and its active time window to pinpoint the exact database operations. With all malicious file and database operations precisely identified, Ancora performs a unified rewind and selective replay recovery. It first reverts the system to a clean snapshot taken before the attack. Then, it selectively re-applies only the legitimate operations to both the file system and the database. This process completely removes the attack’s effects while preserving all concurrent legitimate data. We evaluated Ancora on 10 web applications and 20 CVE-based attack scenarios with concurrency up to 150 connections. Our experiments demonstrate that Ancora achieves 99.9% recovery accuracy with manageable overhead: up to 19.8% response latency increase and 17.8% QPS decrease in worst-case scenarios, and recovery throughput of 110.7 database operations per second and 27.2 affected files per second, effectively preserving legitimate data.
@@ I Introduction
Web applications have become primary targets for cyberattacks [1, 2]. While intrusion detection systems (IDS) offer a first line of defense [3, 4, 5, 6, 7, 8, 9, 10], they are not infallible. Consequently, the ability to precisely recover from successful intrusions to ensure data integrity and service continuity is paramount. Modern applications manage distributed and heterogeneous state, encompassing files scattered across disks, relational and non-relational databases, and external web services [11, 12, 13, 14, 15, 16, 17]. Handling any external request (including attack requests) may involve file operations, database operations, and interactions with external web services.
Industrial solutions typically rely on snapshot-based rollbacks, which indiscriminately discard all post-attack operations, including legitimate user transactions, leading to unacceptable data loss [18, 19, 20, 21]. A more effective paradigm is “surgical” recovery, which precisely analyzes and reverses only malicious changes. The central task of such recovery is to identify all modifications—spanning files, databases, and external web services—caused by a single malicious request. However, achieving this in high-concurrency environments is exceedingly challenging. Interleaved streams of low-level syscalls and database logs from numerous concurrent users create an ambiguous many-to-many mapping, severely complicating the attribution of specific database operations and file modifications to a single originating http request.
@@
  • Taint analysis based on metadata propagation instruments an application’s runtime, such as the language interpreter or middleware, to propagate a unique identifier from each http request to its subsequent operations [12, 13, 14, 15, 16]. This allows for direct correlation. However, this approach is fundamentally limited by its high cost and invasiveness, requiring deep, complex modifications to application frameworks that are impractical in production environments. Furthermore, it is inherently blind to unexpected execution paths created by exploits, such as remote command execution, because these actions bypass the instrumented logic and thus remain untainted and untraceable.
  • Black-box behavior pattern matching avoids instrumentation by using external monitoring and machine learning to infer correlations between high-level http requests and low-level system events [17, 11]. The main limitation of this approach is its diminished accuracy in high-concurrency settings. The dense, interleaved event streams make precise attribution challenging for learning-based models, often resulting in misclassifications. These systems are also typically restricted to monitoring predefined directories for performance reasons, leaving them unable to detect malicious file writes in arbitrary, unpredictable locations.
@@
Inspired by attack investigation techniques from forensics analysis, which trace the causal chain of events forward from an initial point of compromise [22, 23, 24, 25, 26, 27], Ancora begins by using forward analysis to isolate the complete syscall sequence corresponding to a given malicious http request. This sequence alone is sufficient to identify and report all interactions with external web services. However, recovering the internal state requires solving two further challenges at a higher semantic level:
@@
To solve the file challenge, we employ causality tracking based on syscalls [28, 29]. Instead of passively monitoring a few directories, Ancora actively analyzes the request’s syscalls to construct a complete provenance graph. This graph reveals the entire causal chain of process execution triggered by the initial malicious act, making any covert file writes by descendant processes visible, regardless of their location on the file system.
@@
In the spirit of open science, we make our tool available at https://anonymous.4open.science/r/Ancora/.
@@ II-A Provenance Graph
A provenance graph is a directed acyclic graph that captures the causal and informational dependencies among system entities [28]. Nodes in the graph represent subjects, such as processes and threads, and objects, such as files and network sockets. Directed edges represent the information flow and causal relationships between them. Every low-level system event, such as a process u writing to a file v at time t, is abstracted as an edge in the graph [30]. When the object of one event becomes the subject of a subsequent event, a causal chain is formed, providing a theoretical foundation for tracing complex behavioral sequences. Forward analysis on a provenance graph interprets these causal relationships to understand an attack’s impact [29, 31]. It begins from a known malicious source and traverses all subsequent events triggered by it, thereby assessing the full scope of the compromise.
@@ II-B Threat Model
  • The audit logging mechanism and the Ancora system are part of the Trusted Computing Base (TCB). We assume their integrity and availability are maintained during an attack and that they cannot be easily tampered with or disabled. This is a common and necessary assumption for many log-based security analysis systems. [32, 33, 34].
  • The target database can be configured to enable comprehensive logging of modification statements, such as setting log_statement = ’mod’ in PostgreSQL. This assumption, shared with other database recovery research [11, 16, 17], ensures that the raw text of all database operations is captured, providing the necessary data for subsequent analysis and recovery.
  • The database environment possesses common and well-established mechanisms for creating snapshots and performing restorations (e.g., mysqldump for MySQL, pg_dump/pg_restore for PostgreSQL [35, 36, 37]), providing the essential basis for system state regression and the assurance of data consistency.
Scope. Ancora focuses on attacks initiated by external requests that produce observable state changes at the system level. Our model does not directly address attacks that operate purely within the application’s logic without triggering discernible differences in database or file I/O at the syscall level. Furthermore, the following attack vectors are considered out of scope: side-channel attacks, hardware-level exploits (e.g., Spectre), resource-exhaustion denial-of-service attacks (e.g., network floods), passive remote system fingerprinting, various network spoofing attacks, and any threats requiring prior physical access or an established foothold within the internal network. This threat model aligns with the consensus in existing web application recovery research [11, 17, 16], allowing us to concentrate on the core challenge of reversing multi-faceted state corruption caused by external attacks.
@@ III-A System Design Overview
Figure 1: The workflow of Ancora.
The core objective of Ancora is to achieve comprehensive and accurate recovery by precisely identifying and reversing all state changes caused by a malicious request, across both files and databases, while preserving all concurrent legitimate operations. As illustrated in Figure 1, the Ancora workflow proceeds in three stages after a malicious http request is identified by an external IDS or a human analyst:
@@ III-B Request Partition
To address this, the Request Partition module deterministically partitions the commingled stream of system-wide syscall audit logs into discrete behavioral units, each corresponding to a single http request. This module adopts a delimiter-logging approach, inspired by prior work on syscall partitioning [22], which uses lightweight, application-agnostic instrumentation at the web framework level. At critical points in a request’s lifecycle, such as its initiation and context switches, our instrumentation injects special delimiter logs containing a unique request ID into the syscall stream. These delimiters act as markers, delineating the boundaries of different requests’ activities within the continuous log data.
@@ III-C Post-Attack File System Recovery
Execution Dependency Analysis. This stage constructs a complete causal chain based on syscalls to identify the full impact of a malicious request on the file system. Instead of passively monitoring predefined directories, this analysis actively constructs a provenance graph. As shown in Figure 2, the analysis begins with the malicious syscall sequence from the Request Partition module. We treat this sequence as the root of the attack and use process creation syscalls, such as fork and clone, as key events for extending the causal chain. Ancora recursively tracks the activities of these newly spawned processes and threads (and their descendants) throughout the global system log.
@@
2) Write Log for Replay. To enable granular replay for data directories, Ancora captures legitimate write operations’ content. Similar to other systems [11], we use a generic, application-agnostic kernel module that, upon intercepting a write syscall to a data directory, logs the event metadata alongside the full data payload. This logged data is buffered in-memory and written to disk asynchronously by a background thread, minimizing performance impact. These captured payloads form the “write log” for replay, enabling Ancora to re-apply legitimate data changes onto files restored from a baseline.
@@ III-D Post-Attack Database State Recovery
Combining these models reveals the path to precise attribution, as illustrated in Figure 3. When the Malicious Web Request begins at Time_s, it borrows Conn_A. For its entire processing duration (from Time_s to Time_t), it has exclusive use of Conn_A. On the database server, Conn_A is handled exclusively by Database Worker A. Therefore, we can conclude that within the precise time window [Time_s, Time_t], all database operations executed by Database Worker A must have originated from the Malicious Web Request and no other. This spatiotemporal anchor, composed of the network connection tuple (identifying Conn_A) and the processing time window (identifying [Time_s, Time_t]), is the cornerstone of our method. It allows us to deterministically attribute behavior across the two independent, asynchronous execution domains.
@@
The analysis begins with the malicious syscall sequence from the Request Partition module. Ancora parses the network-related syscalls in this sequence, filtering for those targeting the database server’s IP and port. Upon such identification, it extracts the full network tuple. Concurrently, it records the timestamps of the first and last syscalls using this tuple during the request’s lifecycle, defining the time window. This stage outputs one or more of these anchors (Malicious Connection Fragments in Figure 1), serving as input for the next stage.
@@ III-E System State Restoration
The final, recovered system state is achieved by composing the outcomes of the file system and database recovery processes detailed in Section III-C and Section III-D, respectively. Once the Rewind & Selective Replay mechanism has surgically repaired the file system and the database state, the system as a whole reaches the Recovered System State shown in Figure 1. By preserving all concurrent legitimate operations across both stateful components, Ancora ensures business continuity and overcomes the critical data loss problem inherent in coarse-grained, snapshot-only rollback approaches.
@@ IV-A Evaluation Setup
Benchmark Applications and Vulnerabilities. To construct a realistic evaluation environment, we selected 9 open-source web applications and one custom application (see Table I). These applications cover four mainstream languages: PHP, Python, Node.js, Java and three database types - MySQL, PostgreSQL(both the standard version and one with the PostGIS extension), MongoDB - encompassing both SQL and NoSQL models. We reproduced 20 real-world, CVE-based attack scenarios, including high-severity vulnerabilities like SQL injection and RCE. All vulnerable environments, except our custom application, were built using images from the [38] project to ensure realism and reproducibility.
@@
Ground Truth and Baselines. We established ground truth by running tests under single-concurrency conditions and manually labeling the causal relationships between requests and operations. We use Sanare [11] as the baseline for all comparative analyses.
@@ IV-D Case Study
In contrast, Ancora achieved precise recovery in both stages. For the database recovery, Ancora correctly identified only the two malicious writes by tracking the request to its specific database connection (172.18.0.3:50564 -> 172.18.0.2:3306) and worker thread (ID=337278), precisely isolating the operations within their execution window (T_start=<15:00:19.700041883>, T_end=<15:00:19.810219831>). No legitimate data was affected. Similarly, for the file system recovery, Ancora correctly identified the webshell creation. It constructed a full provenance graph from syscalls (Figure 4) that linked the http request through the PHP and forked curl processes to the final file write, allowing the recovery process to correctly delete the webshell without affecting other files.
Figure 4: File Provenance Graph of Case Study.
@@ IV-E Performance Overhead
We stress-tested Ancora using Locust with concurrency up to 150, targeting API endpoints that trigger database and file system writes. In the worst-case scenario, Ancora introduced a 19.8% increase in average response latency and a 17.8% decrease in Queries Per Second (QPS). As shown in Figure 5, the overhead varies with application type and concurrency:
@@
We evaluated scalability by measuring recovery time against an increasing number of entities to restore. As shown in Figure 6, Ancora exhibits excellent scalability. The recovery time for both database and file system restoration scales linearly with the number of replayed operations and repaired files. This demonstrates that the recovery overhead is predictable and manageable even for large-scale incidents.

### v3

@@ Abstract
Modern web application recovery presents a critical dilemma. Coarse-grained snapshot rollbacks cause unacceptable data loss for legitimate users. However, surgically removing an attack’s impact is hindered by a fundamental challenge in high-concurrency environments: it is difficult to attribute the resulting file, database modifications, etc., to a specific attack related request. We present Ancora, a system for precise intrusion recovery in web applications without invasive instrumentation. Our approach first isolates the full sequence of syscalls triggered by a single malicious request. Based on this sequence, Ancora addresses file and database modifications separately. To trace file modifications, it builds a provenance graph that reveals all changes, even those made by exploit-spawned processes. To attribute database operations, a more difficult challenge due to connection pooling, Ancora introduces a novel technique based on a spatiotemporal anchor. This anchor uses the request’s network connection tuple and its active time window to pinpoint the exact database operations. With all malicious file and database operations precisely identified, Ancora performs a unified rewind and selective replay recovery. It first reverts the system to a clean snapshot taken before the attack. Then, it selectively re-applies only the legitimate operations to both the file system and the database. This process completely removes the attack’s effects while preserving all concurrent legitimate data. We evaluated Ancora on 10 web applications and 20 CVE-based attack scenarios with concurrency up to 150 connections. Our experiments demonstrate that Ancora achieves 99.9% recovery accuracy with manageable overhead: up to 19.8% response latency increase and 17.8% QPS decrease in worst-case scenarios, and recovery throughput of 110.7 database operations per second and 27.2 affected files per second, effectively preserving legitimate data.
@@ I Introduction
Web applications have become primary targets for cyberattacks [noauthor_owasp_nodate, shahid2022comparative]. While intrusion detection systems (IDS) offer a first line of defense [cheng2024kairos, zengy2022shadewatcher, wang2022threatrace, wang2020youprovdetector, goyal2024rr-caid, han2020unicorn, xiong2020conan, han2021sigl], they are not infallible. Consequently, the ability to precisely recover from successful intrusions to ensure data integrity and service continuity is paramount. Modern applications manage distributed and heterogeneous state, encompassing files scattered across disks, relational and non-relational databases, and external web services [matos2021sanare, ammann2002recovery, akkucs2010data, chandra2011intrusion, chandra2013asynchronous, nascimento2015shuttle, matos2017rectify]. Handling any external request (including attack requests) may involve file operations, database operations, and interactions with external web services.
Industrial solutions typically rely on snapshot-based rollbacks, which indiscriminately discard all post-attack operations, including legitimate user transactions, leading to unacceptable data loss [veeam, zerto, rubrik, cohesity]. A more effective paradigm is “surgical” recovery, which precisely analyzes and reverses only malicious changes. The central task of such recovery is to identify all modifications—spanning files, databases, and external web services—caused by a single malicious request. However, achieving this in high-concurrency environments is exceedingly challenging. Interleaved streams of low-level syscalls and database logs from numerous concurrent users create an ambiguous many-to-many mapping, severely complicating the attribution of specific database operations and file modifications to a single originating http request.
@@
  • Taint analysis based on metadata propagation instruments an application’s runtime, such as the language interpreter or middleware, to propagate a unique identifier from each http request to its subsequent operations [ammann2002recovery, akkucs2010data, chandra2011intrusion, chandra2013asynchronous, nascimento2015shuttle]. This allows for direct correlation. However, this approach is fundamentally limited by its high cost and invasiveness, requiring deep, complex modifications to application frameworks that are impractical in production environments. Furthermore, it is inherently blind to unexpected execution paths created by exploits, such as remote command execution, because these actions bypass the instrumented logic and thus remain untainted and untraceable.
  • Black-box behavior pattern matching avoids instrumentation by using external monitoring and machine learning to infer correlations between high-level http requests and low-level system events [matos2017rectify, matos2021sanare]. The main limitation of this approach is its diminished accuracy in high-concurrency settings. The dense, interleaved event streams make precise attribution challenging for learning-based models, often resulting in misclassifications. These systems are also typically restricted to monitoring predefined directories for performance reasons, leaving them unable to detect malicious file writes in arbitrary, unpredictable locations.
@@
Inspired by attack investigation techniques from forensics analysis, which trace the causal chain of events forward from an initial point of compromise [wang2023tesec, lee2013high, yu2021alchemist, ma2017mpi, alhanahnah2022autompi, hassan2020omegalog], Ancora begins by using forward analysis to isolate the complete syscall sequence corresponding to a given malicious http request. This sequence alone is sufficient to identify and report all interactions with external web services. However, recovering the internal state requires solving two further challenges at a higher semantic level:
@@
To solve the file challenge, we employ causality tracking based on syscalls [li2021threat, hossain2018dependence]. Instead of passively monitoring a few directories, Ancora actively analyzes the request’s syscalls to construct a complete provenance graph. This graph reveals the entire causal chain of process execution triggered by the initial malicious act, making any covert file writes by descendant processes visible, regardless of their location on the file system.
@@
In the spirit of open science, we make our tool available at https://github.com/ma-b21/Ancora.
@@ II-A Provenance Graph
A provenance graph is a directed acyclic graph that captures the causal and informational dependencies among system entities [li2021threat]. Nodes in the graph represent subjects, such as processes and threads, and objects, such as files and network sockets. Directed edges represent the information flow and causal relationships between them. Every low-level system event, such as a process u writing to a file v at time t, is abstracted as an edge in the graph [xu2022depcomm]. When the object of one event becomes the subject of a subsequent event, a causal chain is formed, providing a theoretical foundation for tracing complex behavioral sequences. Forward analysis on a provenance graph interprets these causal relationships to understand an attack’s impact [hossain2018dependence, zhang2016causality]. It begins from a known malicious source and traverses all subsequent events triggered by it, thereby assessing the full scope of the compromise.
@@ II-B Threat Model
  • The audit logging mechanism and the Ancora system are part of the Trusted Computing Base (TCB). We assume their integrity and availability are maintained during an attack and that they cannot be easily tampered with or disabled. This is a common and necessary assumption for many log-based security analysis systems. [pohly2012hi, bates2015trustworthy, pasquier2017practical].
  • The target database can be configured to enable comprehensive logging of modification statements, such as setting log_statement = ’mod’ in PostgreSQL. This assumption, shared with other database recovery research [matos2021sanare, nascimento2015shuttle, matos2017rectify], ensures that the raw text of all database operations is captured, providing the necessary data for subsequent analysis and recovery.
  • The database environment possesses common and well-established mechanisms for creating snapshots and performing restorations (e.g., mysqldump for MySQL, pg_dump/pg_restore for PostgreSQL [pgdump, pgrestore, mysqldump]), providing the essential basis for system state regression and the assurance of data consistency.
Scope. Ancora focuses on attacks initiated by external requests that produce observable state changes at the system level. Our model does not directly address attacks that operate purely within the application’s logic without triggering discernible differences in database or file I/O at the syscall level. Furthermore, the following attack vectors are considered out of scope: side-channel attacks, hardware-level exploits (e.g., Spectre), resource-exhaustion denial-of-service attacks (e.g., network floods), passive remote system fingerprinting, various network spoofing attacks, and any threats requiring prior physical access or an established foothold within the internal network. This threat model aligns with the consensus in existing web application recovery research [matos2021sanare, matos2017rectify, nascimento2015shuttle], allowing us to concentrate on the core challenge of reversing multi-faceted state corruption caused by external attacks.
@@ III-A System Design Overview
Figure 1: The workflow of Ancora. Starting from a detected malicious http request, Ancora first partitions syscall logs to isolate the request’s syscall sequence, then performs provenance-based file recovery and spatiotemporal-anchor-based database recovery, and finally conducts unified rewind and selective replay to remove all malicious effects while preserving concurrent legitimate data.
The core objective of Ancora is to achieve comprehensive and accurate recovery by precisely identifying and reversing all state changes caused by a malicious request, across both files and databases, while preserving all concurrent legitimate operations. As illustrated in Fig. 1, the Ancora workflow proceeds in three stages after a malicious http request is identified by an external IDS or a human analyst:
@@ III-B Request Partition
To address this, the Request Partition module deterministically partitions the commingled stream of system-wide syscall audit logs into discrete behavioral units, each corresponding to a single http request. This module adopts a delimiter-logging approach, inspired by prior work on syscall partitioning [wang2023tesec], which uses lightweight, application-agnostic instrumentation at the web framework level. At critical points in a request’s lifecycle, such as its initiation and context switches, our instrumentation injects special delimiter logs containing a unique request ID into the syscall stream. These delimiters act as markers, delineating the boundaries of different requests’ activities within the continuous log data.
@@ III-C Post-Attack File System Recovery
Execution Dependency Analysis. This stage constructs a complete causal chain based on syscalls to identify the full impact of a malicious request on the file system. Instead of passively monitoring predefined directories, this analysis actively constructs a provenance graph. As shown in Fig. 2, the analysis begins with the malicious syscall sequence from the Request Partition module. We treat this sequence as the root of the attack and use process creation syscalls, such as fork and clone, as key events for extending the causal chain. Ancora recursively tracks the activities of these newly spawned processes and threads (and their descendants) throughout the global system log.
@@
2) Write Log for Replay. To enable granular replay for data directories, Ancora captures legitimate write operations’ content. Similar to other systems [matos2021sanare], we use a generic, application-agnostic kernel module that, upon intercepting a write syscall to a data directory, logs the event metadata alongside the full data payload. This logged data is buffered in-memory and written to disk asynchronously by a background thread, minimizing performance impact. These captured payloads form the “write log” for replay, enabling Ancora to re-apply legitimate data changes onto files restored from a baseline.
@@ III-D Post-Attack Database State Recovery
Combining these models reveals the path to precise attribution, as illustrated in Fig. 3. When the Malicious Web Request begins at Time_s, it borrows Conn_A. For its entire processing duration (from Time_s to Time_t), it has exclusive use of Conn_A. On the database server, Conn_A is handled exclusively by Database Worker A. Therefore, we can conclude that within the precise time window [Time_s, Time_t], all database operations executed by Database Worker A must have originated from the Malicious Web Request and no other. This spatiotemporal anchor, composed of the network connection tuple (identifying Conn_A) and the processing time window (identifying [Time_s, Time_t]), is the cornerstone of our method. It allows us to deterministically attribute behavior across the two independent, asynchronous execution domains.
@@
The analysis begins with the malicious syscall sequence from the Request Partition module. Ancora parses the network-related syscalls in this sequence, filtering for those targeting the database server’s IP and port. Upon such identification, it extracts the full network tuple. Concurrently, it records the timestamps of the first and last syscalls using this tuple during the request’s lifecycle, defining the time window. This stage outputs one or more of these anchors (Malicious Connection Fragments in Fig. 1), serving as input for the next stage.
@@ III-E System State Restoration
The final, recovered system state is achieved by composing the outcomes of the file system and database recovery processes detailed in Section III-C and Section III-D, respectively. Once the Rewind & Selective Replay mechanism has surgically repaired the file system and the database state, the system as a whole reaches the Recovered System State shown in Fig. 1. By preserving all concurrent legitimate operations across both stateful components, Ancora ensures business continuity and overcomes the critical data loss problem inherent in coarse-grained, snapshot-only rollback approaches.
@@ IV-A Evaluation Setup
Benchmark Applications and Vulnerabilities. To construct a realistic evaluation environment, we selected 9 open-source web applications and one custom application (see Table I). These applications cover four mainstream languages: PHP, Python, Node.js, Java and three database types - MySQL, PostgreSQL(both the standard version and one with the PostGIS extension), MongoDB - encompassing both SQL and NoSQL models. We reproduced 20 real-world, CVE-based attack scenarios, including high-severity vulnerabilities like SQL injection and RCE. All vulnerable environments, except our custom application, were built using images from the [vulhub] project to ensure realism and reproducibility.
@@
Ground Truth and Baselines. We established ground truth by running tests under single-concurrency conditions and manually labeling the causal relationships between requests and operations. We use Sanare [matos2021sanare] as the baseline for all comparative analyses.
@@ IV-D Case Study
In contrast, Ancora achieved precise recovery in both stages. For the database recovery, Ancora correctly identified only the two malicious writes by tracking the request to its specific database connection (172.18.0.3:50564 -> 172.18.0.2:3306) and worker thread (ID=337278), precisely isolating the operations within their execution window (T_start=<15:00:19.700041883>, T_end=<15:00:19.810219831>). No legitimate data was affected. Similarly, for the file system recovery, Ancora correctly identified the webshell creation. It constructed a full provenance graph from syscalls (Fig. 4) that linked the http request through the PHP and forked curl processes to the final file write, allowing the recovery process to correctly delete the webshell without affecting other files.
Figure 4: File Provenance Graph of Case Study. By tracing the causal chain of process spawning and file I/O operations, Ancora precisely links the initial malicious http request to the creation of the hidden webshell.
@@ IV-E Performance Overhead
We stress-tested Ancora using Locust with concurrency up to 150, targeting API endpoints that trigger database and file system writes. In the worst-case scenario, Ancora introduced a 19.8% increase in average response latency and a 17.8% decrease in Queries Per Second (QPS). As shown in Fig. 5, the overhead varies with application type and concurrency:
@@
We evaluated scalability by measuring recovery time against an increasing number of entities to restore. As shown in Fig. 6, Ancora exhibits excellent scalability. The recovery time for both database and file system restoration scales linearly with the number of replayed operations and repaired files. This demonstrates that the recovery overhead is predictable and manageable even for large-scale incidents.